OpenStudy (anonymous):

could anyone help me solve this problem please: sqt x-3 + sqt 2x-2 = 3

OpenStudy (anonymous):

\[\sqrt{x-3} + \sqrt{2x-2} = 3\]\[\sqrt{2x-2} = 3 - \sqrt{x-3}\] Squaring both sides: \[2x-2 = 9 - 6\sqrt{x-3} + x - 3\]\[6\sqrt{x-3} = 8 - x\]Again squaring both sides: \[(6\sqrt{x-3})^{2}=(8-x)^{2}\]

OpenStudy (anonymous):

@viande Welcome to Open Study.

OpenStudy (anonymous):

thank you :) and btw is it necessary for the equations to be transposed?

OpenStudy (anonymous):

not a compulsion.

OpenStudy (anonymous):

If you don't transposed then you will get \[2\sqrt{2x^{2}-8x+6}=8-3x\]